# 61A Bot Report: AI Assistants in CS1 Save Students Homework Time and Reduce Demands on Staff. (Now What?)

### v2

@@ 5.1. Limitations & Threats to Validity
Similarly, differences in course content delivery, staffing, structure, etc., would be expected to impact both labs and homeworks, as lectures and discussion sections for given topics come before labs and homeworks. To our knowledge, no additional homework support was provided in Spring 2024 beyond 61A Bot—no additional hints or support unique to Spring 2024 were offered in lecture or group sessions.
Finally, our study is entirely observational: we had limited levers for randomization, and there could be uncontrolled causes for the effects we observe—for example, it could be that students are using ChatGPT for their homework or projects anyway, despite the prohibition on ChatGPT use and the availability of 61A Bot. ChatGPT was originally released in November 2022, but we observed no similar effect in Spring 2023 compared with prior terms.
@@ Acknowledgements.
This work was made possible by a few generous sources of support: an Inclusion Research Award from Google, and support for 61A-Bot’s use of Azure’s OpenAI API by Microsoft.References

### v3

@@ 5.1. Limitations & Threats to Validity
Similarly, differences in course content delivery, staffing, structure, etc., would be expected to impact both labs and homeworks, as lectures and discussion sections for given topics come before labs and homeworks. No additional homework support was provided in Spring 2024 beyond 61A Bot—no hints or support unique to Spring 2024 were offered in lecture or group sessions.
Finally, our study is entirely observational, and there could be other causes for the effects we observe—perhaps students are using ChatGPT for their homework or projects despite the prohibition on ChatGPT use and the availability of 61A Bot. (ChatGPT was originally released in November 2022, but we observed no similar effect in Spring 2023 compared with prior terms.)
@@ Acknowledgements.
This work was made possible by a few generous sources of support: an Inclusion Research Award from Google, and support for 61A Bot’s use of Azure’s OpenAI API by Microsoft. We would also like to extend our graitiude to our anonymous reviewers for their helpful feedback, as well as to our student participants in this experiment.References